Company 1 and Company 2 are identical firms in all respects except for their capital structure. Company 1 is all equity financed with $800,000 in stock. Company 2 uses both stock and perpetual debt; its stock is worth $400,000 and the interest rate on its debt is 10%. Both firms expect EBIT to be $95,000 and all income will be distributed as dividends. Ignore taxes.

a. Fred owns $30,000 worth of Company 2 stock. What rate of return is he expecting?

b. Show how Fred could generate exactly the same cash flows and rate of return by investing in company 1 and using homemade leverage.

c. Now assume company 1 and company 2 each pay a 20% marginal corporate tax, but Fred pays no taxes. Repeat a) and b). How is the outcome different than in a) and b)? Explain. Which firm would Fred prefer to invest in? Why?

d. Now assume company 1 and company 2 each pay a 20% marginal corporate tax, and Fred pays a 15% tax on dividends. Repeat a) and b). How is the outcome different than in a), b), and c)? Explain. Which firm would Fred prefer to invest in? Why?
Answer #1

1.
Income available for Equity share holders = EBIT - interest on debt = 95000-(0.10*400000) = 55000
Dividend for each $ invested in Company 2 = 55000/400000 = 0.1375
Fred's dividend income = 0.1375*30000 = $4125
Return on his investment = (4125/30000)*100 = 13.75%

2.
To generate exactly the same cash flows in Company 1, sale the shareholding in Company 2 for $30,00. Then borrow amount $30,000 @10% interest. Then use the proceeds of the stock sale and the loan to buy shares in Company 1.
The investor will receive dividends in proportion to the percentage of the company 1’s share they own.
The total dividends received by the shareholder will be: Dividends received = (95,000/800,000)*60,000 Dividends received = $7,125
Therefore, the net total cash flow for the shareholder will be= dividend from stock of company 1 - interest payment on loan = 7215-(0.10*30000) = $4,125
Return on his investment = 4125/30000 = 13.75%

3.
EAT, Company 1 = 95000*(1-0.20) = 76000
EAT, Company 2 = 95000-Interest on loan - taxes = 95000-(0.10*400000) - 0.20*(95000-(0.10*400000)) = $44,000
a. Dividend income on Fred's investment in Company 2 = (44000/400000)*30000 = 3300
Return on his investment = (3300/30000)*100 = 11.00%

b. Net cash flow from Fred's investment in Company 1 = ((76000/800000)*60000)-(0.10*30000) = $2,700
Return on his investment = 2700/30000 = 9.00%

The outcome is different than in (a) and (b) because of the tax savings on interest paid by Company 2. Since, company 2 pays out interest, its net taxable income and therefore the tax liability is lower than that of Company 1, which results in higher per share return for Company 2. Thus, Fred should invest in Company 2 shares.

4.
EAT, Company 1 = 95000*(1-0.20) = 76000
EAT, Company 2 = 95000-Interest on loan - taxes = 95000-(0.10*400000) - 0.20*(95000-(0.10*400000)) = $44,000
a. Dividend income on Fred's investment in Company 2 = ((44000/400000)*30000)*(1-0.15) = 2805
Return on his investment = (2805/30000)*100 = 9.35%
b. Net cash flow from Fred's investment in Company 1:
Dividend Income = ((76000/800000)*60000)*(1-0.15) = 4845
Interest on loan =(0.10*30000) = $3,000
Net Income = 4845-3000 = 1845
Return on his investment = 1845/30000 = 6.15%

The outcome is different from (a) and (b) due to the corporate tax saving effects of interest paid by Company 2. The putcome is different from (c) because of the dividend tax paid by Fred on his dividend income. The dividend income from Company 1 is higher for Fred than for Company 2, resulting in a higher tax liability and a higher percentage reduction in his rate of return on investment. Therefore, Fred should invest in Company 2 shares.